OpenStudy (anonymous):

Solve the proportion 2 times x all over 5 times x plus 6 equals 2 over x.. What is the value of x?

OpenStudy (blurbendy):

2x / (5x + 6) = 2 / x x = -1 and x = 6 both work 2(-1) / (5(-1) + 6) = 2 / (-1) - 2 / 1 = -2 / 1 2(6) / (5(6) + 6) = 2 / (6) 12 / 36 = 1 / 3 1/3 = 1/3

OpenStudy (blurbendy):

if you move the x on the right to the left side of the equation by multiplying by x, you end up with a quadratic equation, hence the two answers

OpenStudy (anonymous):

so are the aswers 1/3?

OpenStudy (blurbendy):

no, they're x = -1, and x = 6

OpenStudy (anonymous):

sorry it took me so long i was trying to read it and understand it

OpenStudy (anonymous):

thank you so much can you help me with other ones?

OpenStudy (blurbendy):

sure

OpenStudy (anonymous):

ok i will send them on this page

OpenStudy (blurbendy):

ok

OpenStudy (anonymous):

What is the exact value of the square root of the quantity 99 over 55.? Simplify if possible.

OpenStudy (anonymous):

\[\sqrt{99/55}\]

OpenStudy (blurbendy):

3 / sqrt[5]

OpenStudy (anonymous):

that is not one of my answer choises do you want the answer choices?

OpenStudy (blurbendy):

yes, please

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

What is the exact value of the square root of the quantity 99 over 55.? Simplify if possible. 3 the square root of 5 all over 5. 3 the square root of 11 all over 5. 11 the square root of 3 all over 5. 5 the square root of 3 all over 5.

OpenStudy (anonymous):

that is the question and answers

OpenStudy (blurbendy):

oh, okay, it's the first one

OpenStudy (blurbendy):

do you see why?

OpenStudy (anonymous):

you devide by 11?

OpenStudy (anonymous):

here is the second one and i have one more and i'm done What is the exact value of the square root of 18 over 3 the square root of 36.? Simplify if possible

OpenStudy (blurbendy):

yes

OpenStudy (blurbendy):

sqrt[2] / 6

OpenStudy (anonymous):

you are so smart i wish i was you

OpenStudy (blurbendy):

the three's get divided out, and you're left with 6 on the bottom

OpenStudy (blurbendy):

lol, ive just practiced this stuff a lot

OpenStudy (anonymous):

oh lucky i barely have time to eat with all the homework i get What is the simplified form of the expression the square root of the quantity 16 times c to the fourth power. - the square root of c squared. + 3the square root of c squared. + the square root of the quantity 9 times c squared.?

OpenStudy (blurbendy):

okay, i got c(4c + 5) or 4c^2 + 5c

OpenStudy (anonymous):

god i'm so lucky to have you helping me

OpenStudy (blurbendy):

no problem

OpenStudy (blurbendy):

anytime!

OpenStudy (anonymous):

how do you give people medals i really wanna give you a medal for all your effort

OpenStudy (blurbendy):

just click the blue button that says Best Response

OpenStudy (blurbendy):

hurray

OpenStudy (anonymous):

thank you so much for all the help and can you help me if i ever need anything in math?

OpenStudy (blurbendy):

of course, just message me. you can know when I'm online by becoming my "fan".

OpenStudy (blurbendy):

Click my name then click "Become a fan"

OpenStudy (anonymous):

already did lol

OpenStudy (blurbendy):

oh okay, cool

OpenStudy (anonymous):

is it ok if you give me your number so i can text you if you ever offline? like not your real number but your textfree or textnow number and app number?

OpenStudy (blurbendy):

I'm not sure what you mean by textfree or textnow number

OpenStudy (anonymous):

it's a texting app and you get a number if you do not want to give me your real number

OpenStudy (blurbendy):

oh, i see. I'll look into it. I'm on here a lot though, so just message me here for now

OpenStudy (anonymous):

ok thank you so much!! ma'am or sir?

OpenStudy (blurbendy):

sir!

OpenStudy (blurbendy):

sir makes me feel old though. just call me blur

OpenStudy (anonymous):

ok blur i love that ps3 game called blur anyways i should let you go bye and thnx once again

OpenStudy (blurbendy):

haha, alright. see you around